# Mini-Monkey: Alleviate the Sawtooth Effect by Multi-Scale Adaptive Cropping

### v1

@@ 4.1 Implementation Details
We use a well-trained InternViT [8], MLP layers, and the InternLLM [4] from InternVL2-2B [7] as the vision encoder, connector and the LLM. Following previous work [8], we use the (448, 448) as the input resolution of InternViT. During the inference, for the detailed Layer, we set the maximum of the sub-image $Max_{num}$ as 24 and set the minimum of the sub-image as 8. For the adaptive layer, we set the maximum of the sub-image as 8 and set the minimum of the sub-image as 1.
Evaluation. Following the previous work [19, 7], we evaluate Mini-Monkey on eleven general multimodal understanding benchmarks, including MathVista testmini [47], SEED Image [31], RealWorldQA [69], AI2D test [27], POPE [34], CCBench [42], MME [16], and HallusionBench [18].
For document understanding, following the previous work [44], we employ two distinct types of metrics to assess the performance of Mini-Monkey. Initially, we leverage the standard metrics provided by the benchmarks to evaluate Mini-Monkey. For this metric, similar to  [7], we utilize benchmarks such as ChartQA [50], DocVQA [53], InfoVQA [52], TextVQA [60], and OCRBench [43]. ChartQA, DocVQA, InfoVQA, and TextVQA are widely used to assess the textual comprehension of models. OCRBench, a more recent benchmark, includes 29 datasets to provide a comprehensive evaluation of the model’s capabilities. Subsequently, we apply the accuracy metric to verify the performance. For this metric, a response from Mini-Monkey that fully captures the ground truth is considered a true positive. Further details on this metric and the used benchmarks can be referenced in [43].
@@
Mini-Monkey (ours) 2B 1881.9 57.5 74.7 75.5 71.3 38.7 86.7 47.3
@@
Mini-Monkey (Ours) 2B 87.4 76.5 60.1 75.7 802
@@
Mini-Monkey-2B (Ours) 66.5 68.4 78.1 49.6 57.9 42.9 70.3 69.9
Table 4: Ablation study of multi-scale adaptive crop strategy. We compare our method with the existing crop strategy and the overlay cropping strategy.
@@
Baseline Dynamic High-Resolution strategy [7] 73.4 784 1876.8 37.9 85.2
Baseline Fixed Size High-Resolution strategy [36] 74.2 772 1824.5 37.6 85.0
@@
Mini-Monkey (Ours) Multi-Scale Adaptive Crop Strategy 75.7 802 1881.9 38.7 86.7
4.2 Comparison to the State of the Art
@@
General Multimodal Understanding. We evaluate Mini-Monkey on general multimodal understanding following  [19, 7]. The results are shown in 1. Mini-Monkey surpasses other 2B-parameter models on 11 benchmarks. On the MathVista and POIE, Mini-Monkey outperforms the previous state-of-the-art method InternVL2-2B by 1% and 1.5%, respectively. On the HallusionBench, Mini-Monkey outperforms MiniCPM-V-2 by 2.6%. These results showcase the ability of Mini-Monkey to handle general multimodal understanding and reasoning tasks.
@@
document understanding. For the first type of metric, the results are presented in Tab. 2. We use a relaxed accuracy measure for ChartQA, ANLS for DocVQA and InfoVQA, and the VQA score for TextVQA. The results indicate that Mini-Monkey achieves state-of-the-art performance among 2B-parameter multimodal large language models. Compared to InternVL2-2B, our method outperforms it by 2.3%, 1.8%, and 1.2% for TextVQA, InfoVQA, and OCRBench, respectively. Due to the small original resolution of ChartVQA, it is less impacted by cropping operations, resulting in a minor improvement from our method. Notably, in the OCRBench, Mini-Monkey even surpasses the 8B-parameter Large Multimodal Model InternVL2-8B and the 9B-parameter Large Multimodal Model GLM4-V by 0.8% and 1.6%, respectively. These results demonstrate the advantages of a multi-scale adaptive crop strategy in enhancing document understanding.
For the accuracy metric, the results are shown in Tab. 3. Mini-Monkey shows an average performance improvement of 14.8% compared to TextMonkey-9B [44], demonstrating the effectiveness of our method. Mini-Monkey also outperforms the state-of-the-art methods on multiple text-related benchmarks. Specifically, Mini-Monkey achieves 49.2% on FUNSD, 70.3% on SROIE, and 69.9% on POIE, outperforming the previous state-of-the-art method by 10.6%, 23.3%, and 42.0%, respectively. These results further indicate the great potential of Mini-Monkey for downstream task applications, such as visual key information extraction.
@@
Model MSAC TextVQA OCRBench MME HallB POPE
@@
MiniCPM-V-2 $\checkmark$ 76.0 (+1.9) 627 (+22) 1819.5 (+10.9) 36.5 (+0.4) 87.1 (+0.8)
InternVL 2 $\times$ 73.4 784 1876.8 37.9 85.2
InternVL 2 $\checkmark$ 75.7 (+2.3) 802 (+18) 1881.9 (+5.1) 38.7 (+0.8) 86.7 (+1.5)
@@
4.3 Ablation Study
In this section, we perform ablation studies on both general multimodal understanding and document understanding benchmarks to validate the effectiveness of our method. We adopt the TextVQA [60], OCRBench [43], HallusionBench [18], MME [16], and RealWorldQA [69] to conduct ablation studies.
Multi-Scale Adaptive Crop Strategy. We conducted ablation studies to investigate the effectiveness of the proposed multi-scale adaptive crop strategy. We compared our method with several alternatives: The dynamic high-resolution strategy [7], which maintains aspect ratios to increase resolution. The fixed-size high-resolution strategy [36], which uses a fixed size to increase resolution. The overlapping cropping strategy, which uses a high-resolution approach but crops with overlay. The multi-scale strategy [58], which introduce a fixed size multi-scale strategy to the MLLM. As presented in Tab. 4, the proposed multi-scale adaptive crop strategy achieved the best results. Our method outperforms the fixed-size multi-scale strategy [58] by 3.3%. Notably, the over-overlay cropping strategy, instead of improving the model’s performance, actually degraded it.
The proposed multi-scale adaptive crop (MSAC) strategy can be seamlessly integrated into crop-based methods. To demonstrate its effectiveness, we incorporated MSAC into various structures of MLLM, such as MiniCPM-V-2 and InternVL 2. As shown in Tab. 5, MSAC consistently enhances the performance across different MLLM structures, thereby validating the effectiveness of our approach.
Scale Compression Mechanism. We compared the proposed Scale Compression Mechanism with the related work FastV [5]. For different methods, we compress the number of visual tokens by 50%. For our method and FastV, we further conduct an experiment with 90% compression. Following FastV’s paper, we set the K in FastV as 2. FastV is a plug-and-play method. Therefore, we conducted this experiment without training the model. As illustrated in Tab. 6, when using 50% compression and 90% compression, our method outperformed FastV by 21.5% and 4.4%, respectively, demonstrating its effectiveness. Both our method and FastV are parameter-free. FastV compresses input tokens, including both visual and textual tokens, within Transformer blocks. In contrast, our method is more targeted by using tokens with high relative information density to compress tokens with low information density.
Figure 3: Qualitative results of Mini-Monkey. (a) Input Image and Ground Truth. (b) The results of using overlapping cropping strategy. OSC represents overlapping cropping strategy. (c) The results of InternVL2-2B and InternVL2-26B. (d) The results of Mini-Monkey.
@@ 4.4 Qualitative Results
In this section, we provide some qualitative results to demonstrate the effectiveness of our method. We first verify that the sawtooth effect is particularly evident in lightweight MLLMs, adopting InternVL2-2B and InternVL2-26B. As shown in Fig. 3(c), InternVL2-26B can answer the questions correctly. However, due to the word ‘classrooms’ and ‘school’ being cropped, InternVL2-2B gives a wrong answer that addresses the text in the bottom left corner of the original image. While Mini-Monkey can overcome this sawtooth effect and provide the correct answer, as presented in Fig. 3(d). Comparing Fig. 3(b) and Fig. 3(d), we can see that the overlapping cropping strategy introduces some hallucinations and cannot answer questions accurately based on the image, whereas our methods can effectively address the sawtooth effect.
5 Discussion
There are some other methods to address this issue. One approach is to use a vision encoder that inherently supports high resolution, such as the Swin-Transformer [45] or SAM [29]. However, the cropping strategy remains the most commonly employed method. This preference stems from the ability to leverage the pre-trained, robust vision encoder CLIP [56]. Why high-resolution encoders are not always used to tackle this problem directly? The reason lies in the resource efficiency of CLIP pre-training. Typically, due to the low-resolution input, CLIP requires fewer resources compared to high-resolution visual encoders. Consequently, CLIP is often chosen as the visual encoder in multimodal large language model (MLLM) systems, with the cropping strategy being used to enhance the input resolution.
@@ 6 Conclusion
In this study, we introduced Mini-Monkey, a lightweight multimodal large language model (MLLM) designed to address the limitations of existing cropping strategies used to enhance MLLMs’ ability to process high-resolution images. Traditional cropping methods often segment objects and connected areas, which limits the recognition of small or irregularly shaped objects and text, a problem particularly pronounced in lightweight MLLMs. To mitigate this, Mini-Monkey employs a multi-scale adaptive crop strategy (MSAC), generating multi-scale representations that allow for the selection of non-segmented objects across different scales. The proposed MSAC can be consistently enhanced across various MLLM architectures. Additionally, we developed a Scale Compression Mechanism (SCM) to reduce the computational overhead of MSAC by compressing image tokens. Our experimental results demonstrate that Mini-Monkey not only achieves leading performance on a variety of general multimodal model understanding tasks but also shows consistent improvements in document understanding tasks. Notably, on the OCRBench benchmark, Mini-Monkey scored 802, surpassing larger 8B-parameter state-of-the-art models like InternVL2-8B. Furthermore, our model and training strategy are exceptionally efficient, requiring only eight RTX 3090 GPUs for training. These results indicate the potential of Mini-Monkey as a powerful and efficient solution for advancing multimodal large language models capabilities in high-resolution image processing.

### v3

@@ 3.1 Implementation Details
We use InternVL2-2B [8] as the Baseline to develop the Mini-Monkey. Following previous work [9], we use the (448, 448) as the input resolution of InternViT. The training datasets used to train the model include DocVQA [52], ChartQA [50], DVQA [28], AI2D [29], GeoQA+ [5], and LLaVA-150K (zh) [42]. We use the AdamW [47] as the optimizer. The base learning rate is 4e-8.
Evaluation. Following the previous work [21, 8], we evaluate Mini-Monkey on eleven general multimodal understanding benchmarks, including MathVista testmini [49], SEED Image [32], RealWorldQA [64], AI2D test [29], POPE [35], CCBench [43], MME [18], and HallusionBench [20]. For document understanding, following the previous work [45], we employ two distinct types of metrics to verify the performance of Mini-Monkey. Initially, we leverage the standard metrics provided by the benchmarks to evaluate Mini-Monkey. We utilize benchmarks such as ChartQA [50], DocVQA [52], InfoVQA [51], TextVQA [56], STVQA [3], FUNSD [27], SROIE [26], POIE [30] and OCRBench [44]. We also apply the accuracy metric to verify the performance. For this metric, a response from Mini-Monkey that fully captures the ground truth is considered a true positive. Further details on this metric and the used benchmarks can be referenced in appendix LABEL:more_implementation.
@@
Mini-Monkey (ours) 2B 1884.2 57.9 74.8 75.5 71.3 38.8 88.0 47.3
@@
Mini-Monkey (Ours) 2B 87.4 76.5 60.1 76.0 806
@@
InternVL2-2B [8] 65.6 66.2 76.7 46.8 67.6 42.0 68.0 66.8
Mini-Monkey-2B (Ours) 67.2 68.8 78.4 50.0 67.3 43.2 70.5 71.2
Table 4: Ablation study of Complementary Image Pyramid. We compare our method with the existing cropping strategy and the overlay cropping strategy.
@@
Baseline Dynamic high-res Strategy [8] 73.4 784 1876.8 37.9 85.2
Baseline Fixed Size high-res Strategy [36] 74.2 772 1824.5 37.6 85.0
@@
Mini-Monkey (Ours) Complementary Image Pyramid 76.0 806 1884.2 38.8 88.0
3.2 Comparison with State-of-the-art Methods
@@
General Multimodal Understanding. We evaluate Mini-Monkey on general multimodal understanding following  [21, 8]. The results are shown in Tab. 1. Mini-Monkey surpasses other 2B-parameter models on 8 benchmarks. On the MathVista and POPE, Mini-Monkey outperforms the previous state-of-the-art method InternVL2-2B by 1% and 2.8%, respectively. On the HallusionBench, Mini-Monkey outperforms MiniCPM-V-2 by 2.7%. These results showcase the ability of Mini-Monkey to handle general multimodal understanding and reasoning tasks.
@@
Document Understanding. For the first type of metric, the results are presented in Tab. 2. We use a relaxed accuracy measure for ChartQA, ANLS for DocVQA and InfoVQA, and the VQA score for TextVQA. The results indicate that Mini-Monkey achieves state-of-the-art performance among 2B-parameter multimodal large language models. Compared to InternVL2-2B, our method outperforms it by 2.6%, 1.2%, and 22 for TextVQA, InfoVQA, and OCRBench, respectively. Due to the small original resolution of ChartQA, it is less impacted by cropping operations, resulting in a minor improvement from our method. Notably, in the OCRBench, Mini-Monkey even surpasses the 8B-parameter Large Multimodal Model InternVL2-8B and the 9B-parameter Large Multimodal Model GLM4-V by 12 and 20, respectively. These results demonstrate the advantages of a Complementary Image Pyramid in enhancing document understanding. For the accuracy metric, the results are shown in Tab. 3. Mini-Monkey shows an average performance improvement of 15.6% compared to TextMonkey-9B [45], demonstrating the effectiveness of our method. Mini-Monkey also outperforms the state-of-the-art methods on multiple text-related benchmarks. Specifically, Mini-Monkey outperforms the InternVL2-2B by 2.6%, 3.2%, and 4.4% on TextVQA, InfoVQA, and POIE, respectively. These results further indicate the great potential of Mini-Monkey for document understanding tasks.
3.3 Ablation Study
In this section, we perform ablation studies on both general multimodal understanding and document understanding benchmarks to validate the effectiveness of our method. We adopt the TextVQA [56], OCRBench [44], HallusionBench [20], MME [18], and POPE [35] to conduct ablation studies.
Complementary Image Pyramid. We conducted ablation studies to investigate the effectiveness of the CIP. We compared our method with several alternatives: The dynamic high-res strategy [8], which maintains aspect ratios to increase resolution. The fixed-size high-res strategy [36], which uses a fixed size to increase resolution. The overlapping cropping strategy uses a high-res approach but crops with overlay. The multi-scale strategy [55], which introduces a multi-scale strategy to the MLLM. As presented in Tab. 4, the proposed CIP achieved the best results. Our method outperforms the previous multi-scale strategy [55] on both general multimodal understanding and document understanding. Notably, the over-overlay cropping strategy, instead of improving the model’s performance, actually degraded it.
Various Model Capacity. We performed ablation studies to assess the impact of CIP on models with varying model capacity. As illustrated in Table 6, CIP consistently improves the performance of varying model capacity.
Different Usage Configurations. To further validate the improvements introduced by CIP, we performed experiments on usage configurations: a training-free configuration and a fine-tuning configuration. As shown in Tab. 5, CIP demonstrates improvements in performance even when applied without training. The performance can be further improved with fine-tuning. Additionally, we surprisingly find that CIP can even facilitate the model fine-tuning process. As presented in the second line in Tab. 5, direct fine-tuning of the Baseline model not only failed to improve performance but, in some cases, led to a decline. Conversely, incorporating CIP during the fine-tuning of the Baseline resulted in substantial improvements in both general multimodal understanding and document understanding, as evidenced in the fourth line of Tab. 5.
Incorporating CIP to various MLLMs. The proposed complementary image pyramid (CIP) can be seamlessly integrated into crop-based methods. To demonstrate its effectiveness, we incorporated CIP into various structures of MLLM, such as MiniCPM-V-2 [66], InternVL 2 [8], LLaVA-OV [31]. As shown in Tab. 6, CIP consistently enhances the performance across different MLLM structures, thereby validating the effectiveness of our approach.
The Number of Sub-Images. To investigate whether the performance enhancement is attributed to an increase in the number of sub-images, we performed an experiment by incrementally raising the sub-image count for the Baseline. The findings, summarized in Table 8, indicate that increasing the number of sub-images does not lead to better performance; instead, it may result in a decline. In contrast, CIP can effectively improve the performance of the model. These results further demonstrate the effectiveness of the CIP.
Table 5: Exploring different usage configurations of CIP. Train represents fine-tuning the model.
Model CIP Train TextVQA OCRBench MME HallB POPE
InternVL2-2B $\times$ $\times$ 73.4 784 1876.8 37.9 85.2
InternVL2-2B $\times$ $\checkmark$ 73.3 (-0.1) 787 (+3) 1858.3 (-18.5) 37.3 (-0.6) 85.3 (+0.1)
InternVL2-2B $\checkmark$ $\times$ 75.2 (+1.8) 800 (+17) 1881.9 (+5.1) 38.7 (+0.8) 86.7 (+1.5)
InternVL2-2B $\checkmark$ $\checkmark$ 76.0 (+2.6) 806 (+22) 1884.2 (+7.4) 38.8 (+0.9) 88.0 (+2.8)
@@ 3.3 Ablation Study
Model CIP TextVQA OCRBench MME HallB POPE
@@
MiniCPM-V-2-2.8B $\checkmark$ 76.0 (+1.9) 627 (+22) 1819.5 (+10.9) 36.5 (+0.4) 87.1 (+0.8)
LLaVA-OV-0.5B $\times$ 65.3 577 1478.0 28.1 86.7
LLaVA-OV-0.5B $\checkmark$ 66.2 (+0.9) 600 (+23) 1482.6 (+4.6) 28.8 (+0.7) 87.7 (+1.0)
[…]
InternVL2-1B $\checkmark$ 72.3 (+1.8) 772 (+18) 1801.5 (+7.1) 34.3 (+0.9) 85.7 (+0.8)
InternVL2-2B $\times$ 73.4 784 1876.8 37.9 85.2
InternVL2-2B $\checkmark$ 76.0 (+2.6) 806 (+22) 1884.2 (+7.4) 38.8 (+0.9) 88.0 (+2.8)
[…]
InternVL2-8B $\checkmark$ 79.3 (+1.9) 809 (+15) 2226.4 (+16.1) 45.4 (+0.4) 84.8 (+0.6)
@@
Scale Compression Mechanism. We compared the proposed Scale Compression Mechanism with the related work FastV [6]. For different methods, we compress the number of visual tokens by 50%. For our method and FastV, we further conduct an experiment with 90% compression. Following FastV’s paper, we set the K in FastV as 2. As illustrated in Tab. 7, when using 50% compression and 90% compression, our method outperformed FastV by 21.5% and 4.4%, respectively, demonstrating its effectiveness. FastV compresses input tokens, including both visual and textual tokens, within Transformer blocks. In contrast, our method works in conjunction with the CIP and more targeted by using tokens with high relative information density to compress tokens with low information density.
Table 8: Ablation study on the number of sub-images. The number denotes the sub-image count.
Model Number TextVQA OCRBench MME HallB POPE
Baseline 18 74.2 782 1851.7 37.0 85.8
Baseline 24 74.4 783 1857.6 36.9 85.8
Baseline 32 74.3 782 1845.0 36.9 85.9
Baseline 48 74.0 767 1841.6 36.2 85.7
CIP 32 76.0 806 1884.2 38.8 88.0
Figure 3: Qualitative results of Mini-Monkey. (a) Input Image and Ground Truth. (b) The results of using overlapping cropping strategy. OSC represents the overlapping cropping strategy. (c) The results of InternVL2-2B and InternVL2-26B. (d) The results of Mini-Monkey.
@@ 3.4 Qualitative Results
In this section, we provide some qualitative results to demonstrate the effectiveness of our method. First, we verify that the semantic sawtooth effect is particularly evident in lightweight MLLMs, which adopt InternVL2-2B and InternVL2-26B. As shown in Fig. 3(c), InternVL2-26B can answer the questions correctly. However, due to the word ‘classrooms’ and ‘school’ being cropped, InternVL2-2B gives a wrong answer that addresses the text in the bottom left corner of the original image. Mini-Monkey can overcome this semantic sawtooth effect and provide the correct answer, as presented in Fig. 3(d). Comparing Fig. 3(b) and Fig. 3(d), we can see that the overlapping cropping strategy introduces some hallucinations and cannot answer questions accurately based on the image, whereas our methods can effectively address the semantic sawtooth effect.
@@ 4 Conclusion
In this study, we introduce a Complementary Image Pyramid (CIP) designed to alleviate the semantic sawtooth effect for MLLMs, thereby enhancing their capability to process high-resolution images effectively. CIP is plug-and-play and can be seamlessly integrated into various multimodal large language models at a low cost. We demonstrate the effectiveness of the proposed CIP across diverse architectures, various parameters, and different usage configurations, leading to consistent performance improvements. Besides, we present a Scale Compression Mechanism (SCM) to compress the visual tokens for computational efficiency. CIP not only enhances the general multimodal understanding performance but also shows consistent improvements in document understanding tasks. Furthermore, our experimental results demonstrate that 2B-parameter MLLM equipped with CIP even surpasses larger 8B-parameter state-of-the-art models like InternVL2-8B on the OCRBench. Limitations: To ensure the seamless application of our method across various architectures, we adopt an image-centric approach to construct an image pyramid without introducing additional parameters. In future work, we will explore the use of trainable Feature Pyramid Network (FPN) for MLLMs, aiming to more efficiently leverage multi-scale features.